OpenStudy (anonymous):

A person runs 3 miles in a half-hour. What is her speed in miles per hour? 0.083 mile per hour 1.5 miles per hour 3 miles per hour 6 miles per hour

OpenStudy (whpalmer4):

To find a rate, divide the amount done by the time it took to do. \[\frac{3 \text{ miles}}{\frac{1}{2} \text{ hour}} = \text{______} \frac{\text{miles}}{\text{hour}}\]

OpenStudy (anonymous):

so 6 miles per hour

OpenStudy (whpalmer4):

Exactly!